OpenStudy (cwtan):

A metal detector has a current of 60mA What is the number of free electron that flow across a cross section of the conductor in 40second?

OpenStudy (anonymous):

1 * 10^19 (apprx)

OpenStudy (ash2326):

Current= \(60mA\) or \(60\times 10^{-3}A\) We know current is charge divided by time ( in seconds) so \[60 \times 10^{-3}= \frac{Q}{40}\] So \[Q=40 \times 60 \times 10^{-3}=2.4 Coulombs\] Now no. of electrons flowing= total charge/(charge of an electron) \[\text{No. of electrons}=\frac{2.4}{1.6 \times 10^{-19}}\] I think you can solve now

OpenStudy (cwtan):

Thanks~ solved